# Blind Inverse Game Theory: Jointly Decoding Rewards and Rationality in Entropy-Regularized Competitive Games

### v1

@@ 4.4 Theoretical Guarantees
Theorem 2 (Finite Sample Bounds for Blind-IGT).
@@
$\|\hat{\theta}-\theta^{*}\|_{2}$ $\leq C_{\theta}\cdot\sqrt{\frac{(m+n)\log(1/\delta)}{N}},$ (8)
$|\hat{\tau}-\tau^{*}|$ $\leq C_{\tau}\cdot\sqrt{\frac{(m+n)\log(1/\delta)}{N}}.$ (9)
@@
Proof Sketch (Appendix C.3).
The proof utilizes concentration inequalities (Lemma 1) to bound the perturbation of the empirical system ($\hat{X},\hat{y}$) (Lemma 2). We then apply standard perturbation theory for least squares (Higham2002) to bound the error in the intermediate solution $\hat{\theta}_{LS}$. Finally, a stability analysis of the normalization step (Lemma 3) confirms the optimal rate for both $\hat{\theta}$ and $\hat{\tau}$. ∎
@@ 4.5 Robustness to Misspecified Normalization
We analyze the behavior of the NLS estimator when the normalization constant $C$ is misspecified, validating Remark 1.
@@
$\hat{\tau}=\frac{C_{\text{assumed}}}{\|\hat{\theta}_{LS}\|_{2}}\approx\frac{C_{\text{assumed}}}{C_{\text{true}}/\tau^{*}}=\tau^{*}\cdot\left(\frac{C_{\text{assumed}}}{C_{\text{true}}}\right).$ (10)
@@ 4.6 Partial Identification and Confidence Sets
When the Rank Condition of Theorem 1 fails (i.e., $\text{rank}(X^{*})<d$), the parameters $(\theta^{*},\tau^{*})$ are not uniquely identifiable, even with the normalization constraint. In this scenario, we shift our focus from point estimation to partial identification, aiming to characterize the set of all parameters consistent with the observed QRE.
@@
$\Theta^{*}=\{(\theta,\tau)\in\mathbb{R}^{d}\times\mathbb{R}_{>0}:X^{*}\theta=\tau y^{*},\|\theta\|_{2}=C\}.$ (11)
When the Rank Condition fails, this set contains multiple elements. Unlike the known-$\tau$ case where the feasible set is an affine subspace (Liao2025), here $\Theta^{*}$ is the intersection of a linear subspace (defined by the bilinear constraint) and a non-convex constraint (the normalization sphere).
@@
$\hat{\Theta}_{N}=\{(\theta,\tau):\|\hat{X}\theta-\tau\hat{y}\|_{2}^{2}\leq\kappa_{N},\|\theta\|_{2}=C,\tau>0\}.$ (12)
@@
Assumption 4 (Bounded Temperature).
@@
Proposition 1 (Coverage of the Confidence Set).
Under Assumptions 1-3 and Assumption 4. Let $\epsilon_{N}=\mathcal{O}(\sqrt{(m+n)\log(1/\delta)/N})$ be the policy estimation error rate (Lemma 1). If we choose the threshold $\kappa_{N}$ such that
@@
where $C_{X},C_{Y}$ are the perturbation constants (Lemma 2), then with probability at least $1-\delta$:
@@
Proof Sketch (Appendix C.4).
We analyze the empirical residual $\|\hat{X}\theta^{*}-\tau^{*}\hat{y}\|_{2}$ for any $(\theta^{*},\tau^{*})\in\Theta^{*}$. Since $X^{*}\theta^{*}=\tau^{*}y^{*}$, the residual depends only on the estimation errors $(\hat{X}-X^{*})$ and $(\hat{y}-y^{*})$. By applying the triangle inequality and leveraging the perturbation bounds (Lemma 2) along with Assumption 4, we establish an upper bound on the residual, which dictates the choice of $\kappa_{N}$. ∎
Proposition 1 guarantees that the constructed confidence set $\hat{\Theta}_{N}$ reliably captures all feasible reward and temperature parameters consistent with the observed behavior, providing a comprehensive solution to the Blind-IGT problem even when strong identifiability fails.
@@ 5.1 Problem Formulation and Identifiability
In the Markov game setting, we aim to recover the reward function $r$ and the temperature $\tau^{*}$ from observed trajectories. We adopt an assumption common in the theoretical analysis of RL with function approximation (Jin2020).
Assumption 5 (Linear Q-Functions).
@@
We adopt Assumption 5 as a necessary starting point for rigorous theoretical analysis in dynamic settings. While our primary theoretical analysis (Theorem 3) assumes the transition dynamics $P$ are known for analytical tractability, Algorithm 2 is readily applicable when $P$ is estimated from data, as demonstrated empirically in Section 6.4.
At each state $s$, the QRE policies $(\mu^{*}(s),\nu^{*}(s))$ satisfy the linearized constraints analogous to (4). We define the local matrices $X(s)$ and vectors $y(s)$. The constraints are:
$X(s;\mu^{*},\nu^{*})\theta^{*}=\tau^{*}\cdot y(s;\mu^{*},\nu^{*}),\quad\forall s\in\mathcal{S}.$ (13)
@@
Proposition 2 (Identifiability in Markov Games).
Under Assumption 5 (with $R=1$), the pair $(\theta^{*},\tau^{*})$ is uniquely identifiable if the aggregated matrix $X^{*}$ has full column rank (rank $d$) and $y^{*}\neq 0$.
@@ 5.2 Estimation Algorithm and Guarantees
Algorithm 2 Model-Based Blind-IGT for Markov Games (Linear Q)
1:Input: Trajectories $\mathcal{D}$ (including $s^{\prime}$), features $\phi$, Normalization $R$.
[…]
3:Estimate Transition Dynamics $\hat{P}$ from $\mathcal{D}$ (e.g., MLE).
4:Construct aggregated matrices $\hat{X}$ and vectors $\hat{y}$.
[…]
6:Recover Rewards:
7:Compute $\hat{Q}(s,a,b)=\langle\phi(s,a,b),\hat{\theta}\rangle$.
8:Compute $\hat{V}(s)$ using $\hat{Q},\hat{\mu},\hat{\nu},\hat{\tau}$.
9:Recover rewards (using $\hat{P}$ or known $P$):
[…]
11:Output: Rewards $\hat{r}$, Temperature $\hat{\tau}$.
We provide theoretical guarantees for Algorithm 2 assuming known dynamics, requiring sufficient coverage of the state space.
Assumption 6 (Coverage).
@@
Theorem 3 (Sample Complexity for MGs).
Under Assumptions 3 (extended to MGs), 5 (with $R=1$), and 6, and assuming the identifiability conditions of Proposition 2 hold and $P$ is known. Let $K$ be the total number of state-action samples. If $K$ is sufficiently large, then with probability at least $1-\delta$:
$\|\hat{\theta}-\theta^{*}\|_{2}$ $=\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}|(m+n)\log(|\mathcal{S}|/\delta)}{K}}\right),$ (14)
$|\hat{\tau}-\tau^{*}|$ $=\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}|(m+n)\log(|\mathcal{S}|/\delta)}{K}}\right),$ (15)
$\|\hat{r}-r^{*}\|_{\infty}$ $=\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}|(m+n)\log(|\mathcal{S}|/\delta)}{K}}\right).$ (16)
@@
Proof Sketch (Appendix D.2).
We first establish uniform convergence of policy estimates across states. The bounds for $\hat{\theta}$ and $\hat{\tau}$ follow the analysis of Theorem 2 applied to the aggregated system. The key challenge is bounding the reward recovery error. This is achieved by establishing the Lipschitz continuity of the V-function with respect to its components ($Q,\mu,\nu,\tau$), thereby controlling the error propagation through the Bellman equation. ∎
Remark 2 (Unknown Dynamics).
Theorem 3 provides guarantees assuming access to the true dynamics $P$. When $P$ is unknown and estimated as $\hat{P}$ (Algorithm 2), the error analysis must account for the uncertainty in $\hat{P}$. However, we demonstrate empirically in Section 6.4 that the optimal $\mathcal{O}(K^{-1/2})$ rate is robustly achieved when estimating $P$ from the same trajectories.
@@ 6 EXPERIMENTS
We conduct a comprehensive empirical evaluation of the Blind-IGT framework and the NLS algorithm. Our experiments aim to: (1) validate the theoretical convergence rates; (2) demonstrate the necessity of Blind-IGT when the temperature is unknown; and (3) assess robustness to model misspecification and unknown dynamics. (See Appendix E for detailed setup).
@@ 6.1 Experimental Setup
Markov Games. We simulate a zero-sum Markov Game with $|\mathcal{S}|=8$ states, $m=n=5$ actions, and discount factor $\gamma=0.9$. We adopt the Linear Q-function parameterization (Assumption 5) with $d=6$. The true temperature is set to $\tau^{*}=1.5$.
@@ 6.2 Validation of Convergence Rates
We validate the finite-sample guarantees established in Theorem 2 and Theorem 3.
@@
Markov Games Results. Figure 3 presents the results for the dynamic setting (assuming known P). The empirical slopes for the Q-parameters ($\hat{\theta}$) and the recovered reward function ($\hat{r}$) are -0.52, and -0.54 for temperature ($\hat{\tau}$). This strongly validates Theorem 3.
Figure 3: Convergence analysis in Markov Games (Theorem 3). The empirical slopes for Q-parameters ($\theta$), temperature ($\tau$), and the recovered reward function ($r$) closely match the theoretical $\mathcal{O}(K^{-1/2})$ rate.
Table 1: Comparison of Blind-IGT (NLS) vs. Standard IGT Liao2025 with known/misspecified temperature ($\tau^{*}=2.0$). Blind-IGT achieves near-oracle performance without prior knowledge of $\tau^{*}$.
@@ 6.3 The Necessity of Blind-IGT
We investigate the performance when the transition dynamics $P$ are unknown and must be estimated from data, relaxing the assumption made for Theorem 3.
We use the Markov Game setup ($|\mathcal{S}|=8,m=n=5,d=6$). We apply Algorithm 2, estimating the transition dynamics $\hat{P}$ via Maximum Likelihood Estimation (with Laplace smoothing) from the observed trajectories. We compare this approach (Estimated-P) against an Oracle that uses the true dynamics (Known-P) during the reward recovery step.
@@ 7 CONCLUSION
This work establishes a new statistical foundation for Inverse Game Theory by removing the restrictive assumption of known agent rationality. By rigorously addressing the bilinear coupling between rewards ($\theta$) and temperature ($\tau$), we resolved the fundamental scale ambiguity that previously precluded identification in entropy-regularized competitive games. Our analysis provides a complete characterization of identifiability, spanning unique point recovery via the statistically optimal NLS estimator and partial identification through rigorous confidence sets. The Blind-IGT framework demonstrates significant robustness, notably achieving the optimal $\mathcal{O}(N^{-1/2})$ convergence rates in dynamic Markov games even when transition dynamics must be estimated from data. This foundational analysis provides the essential theoretical tools for extending statistically rigorous IGT to complex environments involving non-linear function approximation and heterogeneous agent populations.
Supplementary Materials
APPENDICES
Appendix A Notations
@@
$r$ Reward function (Markov Games).
$\tau$ Temperature parameter (entropy regularization weight).
@@
$\gamma$ Discount factor (Markov Games).
$P$ Transition dynamics (Markov Games).
@@
$d$ Dimension of the feature space.
$\phi$ Feature map. $L=\sup\|\phi\|_{2}$.
$\theta^{*}$ True reward (or Q-function) parameters.
$C,R$ Normalization constants for $\|\theta^{*}\|_{2}$.
@@
$\xi$ Soft-Min Gap (lower bound on QRE probabilities).
$C_{\min}$ Coverage lower bound (Markov Games).
System Matrices (Blind-IGT Formulation)
@@
$X(\mu,\nu)$ Aggregated feature matrix $A(\nu)^{\top},B(\mu)^{\top}]^{\top}$.
$y(\mu,\nu)$ Aggregated log-ratio vector $c(\mu)^{\top},d(\nu)^{\top}]^{\top}$.
@@
$\hat{\theta}_{LS}$ Intermediate Least Squares solution (estimate of $\theta^{*}/\tau^{*}$).
$\theta^{\prime*}$ True unnormalized parameters $\theta^{*}/\tau^{*}$.
Appendix B Background and Derivations

### v2

@@ 4.4 Theoretical Guarantees
Theorem 4.6 (Finite Sample Bounds for Blind-IGT).
@@
$\|\hat{\theta}-\theta^{*}\|_{2}$ $\leq C_{\theta}\cdot\sqrt{\frac{(m+n)\log(1/\delta)}{N}},$ (4.2)
$|\hat{\tau}-\tau^{*}|$ $\leq C_{\tau}\cdot\sqrt{\frac{(m+n)\log(1/\delta)}{N}}.$ (4.3)
@@
Proof Sketch (Appendix D.3).
The proof utilizes concentration inequalities (Lemma D.1) to bound the perturbation of the empirical system ($\hat{X},\hat{y}$) (Lemma D.2). We then apply standard perturbation theory for least squares [Higham, 2002] to bound the error in the intermediate solution $\hat{\theta}_{LS}$. Finally, a stability analysis of the normalization step (Lemma D.3) confirms the optimal rate for both $\hat{\theta}$ and $\hat{\tau}$. ∎
@@ 4.5 Robustness to Misspecified Normalization
We analyze the behavior of the NLS estimator when the normalization constant $C$ is misspecified, validating Remark 4.3.
@@
$\hat{\tau}=\frac{C_{\text{assumed}}}{\|\hat{\theta}_{LS}\|_{2}}\approx\frac{C_{\text{assumed}}}{C_{\text{true}}/\tau^{*}}=\tau^{*}\cdot\left(\frac{C_{\text{assumed}}}{C_{\text{true}}}\right).$ (4.4)
@@ 4.6 Partial Identification and Confidence Sets
When the Rank Condition of Theorem 4.4 fails (ie., $\text{rank}(X^{*})<d$), the parameters $(\theta^{*},\tau^{*})$ are not uniquely identifiable, even with the normalization constraint. In this scenario, we shift our focus from point estimation to partial identification, aiming to characterize the set of all parameters consistent with the observed QRE.
@@
$\Theta^{*}=\{(\theta,\tau)\in\mathbb{R}^{d}\times\mathbb{R}_{>0}:X^{*}\theta=\tau y^{*},\|\theta\|_{2}=C\}.$ (4.5)
When the Rank Condition fails, this set contains multiple elements. Unlike the known-$\tau$ case where the feasible set is an affine subspace [Liao et al., 2025], here $\Theta^{*}$ is the intersection of a linear subspace (defined by the bilinear constraint) and a non-convex constraint (the normalization sphere).
@@
$\hat{\Theta}_{N}=\{(\theta,\tau):\|\hat{X}\theta-\tau\hat{y}\|_{2}^{2}\leq\kappa_{N},\|\theta\|_{2}=C,\tau>0\}.$ (4.6)
@@
Assumption 4.7 (Bounded Temperature).
@@
Proposition 4.8 (Coverage of the Confidence Set).
Under Assumptions 4.1-4.5 and Assumption 4.7. Let $\epsilon_{N}=\mathcal{O}(\sqrt{(m+n)\log(1/\delta)/N})$ be the policy estimation error rate (Lemma D.1). If we choose the threshold $\kappa_{N}$ such that
@@
where $C_{X},C_{Y}$ are the perturbation constants (Lemma D.2), then with probability at least $1-\delta$:
@@
Proof Sketch (Appendix D.4).
We analyze the empirical residual $\|\hat{X}\theta^{*}-\tau^{*}\hat{y}\|_{2}$ for any $(\theta^{*},\tau^{*})\in\Theta^{*}$. Since $X^{*}\theta^{*}=\tau^{*}y^{*}$, the residual depends only on the estimation errors $(\hat{X}-X^{*})$ and $(\hat{y}-y^{*})$. By applying the triangle inequality and leveraging the perturbation bounds (Lemma D.2) along with Assumption 4.7, we establish an upper bound on the residual, which dictates the choice of $\kappa_{N}$. ∎
Proposition 4.8 guarantees that the constructed confidence set $\hat{\Theta}_{N}$ reliably captures all feasible reward and temperature parameters consistent with the observed behavior, providing a comprehensive solution to the Blind-IGT problem even when strong identifiability fails.
@@ 5.1 Problem Formulation and Identifiability
In the Markov game setting, we aim to recover the reward function $r$ and the temperature $\tau^{*}$ from observed trajectories. We adopt an assumption common in the theoretical analysis of RL with function approximation [Jin et al., 2020].
Assumption 5.1 (Linear Q-Functions).
@@
We adopt Assumption 5.1 as a necessary starting point for rigorous theoretical analysis in dynamic settings. While our primary theoretical analysis (Theorem 5.4) assumes the transition dynamics $P$ are known for analytical tractability, our algorithm is readily applicable when $P$ is estimated from data, as demonstrated empirically in Section 6.4.
At each state $s$, the QRE policies $(\mu^{*}(s),\nu^{*}(s))$ satisfy the linearized constraints analogous to (3.4). We define the local matrices $X(s)$ and vectors $y(s)$. The constraints are:
$X(s;\mu^{*},\nu^{*})\theta^{*}=\tau^{*}\cdot y(s;\mu^{*},\nu^{*}),\quad\forall s\in\mathcal{S}.$ (5.1)
@@
Proposition 5.2 (Identifiability in Markov Games).
Under Assumption 5.1 (with $R=1$), the pair $(\theta^{*},\tau^{*})$ is uniquely identifiable if the aggregated matrix $X^{*}$ has full column rank (rank $d$) and $y^{*}\neq 0$.
@@ 5.2 Estimation Algorithm and Guarantees
We provide theoretical guarantees assuming known dynamics, requiring sufficient coverage of the state space.
Assumption 5.3 (Coverage).
@@
Theorem 5.4 (Sample Complexity for MGs).
Under Assumptions 4.5 (extended to MGs), 5.1 (with $R=1$), and 5.3, and assuming the identifiability conditions of Proposition 5.2 hold and $P$ is known. Let $K$ be the total number of state-action samples. If $K$ is sufficiently large, then with probability at least $1-\delta$:
$\|\hat{\theta}-\theta^{*}\|_{2}$ $=\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}|(m+n)\log(|\mathcal{S}|/\delta)}{K}}\right),$ (5.2)
$|\hat{\tau}-\tau^{*}|$ $=\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}|(m+n)\log(|\mathcal{S}|/\delta)}{K}}\right),$ (5.3)
$\|\hat{r}-r^{*}\|_{\infty}$ $=\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}|(m+n)\log(|\mathcal{S}|/\delta)}{K}}\right).$ (5.4)
@@
Proof Sketch (Appendix E.2).
We first establish uniform convergence of policy estimates across states. The bounds for $\hat{\theta}$ and $\hat{\tau}$ follow the analysis of Theorem 4.6 applied to the aggregated system. The key challenge is bounding the reward recovery error. This is achieved by establishing the Lipschitz continuity of the V-function with respect to its components ($Q,\mu,\nu,\tau$), thereby controlling the error propagation through the Bellman equation. ∎
Remark 5.5 (Unknown Dynamics).
Theorem 5.4 provides guarantees assuming access to the true dynamics $P$. When $P$ is unknown and estimated as $\hat{P}$, the error analysis must account for the uncertainty in $\hat{P}$. However, we demonstrate empirically in Section 6.4 that the optimal $\mathcal{O}(K^{-1/2})$ rate is robustly achieved when estimating $P$ from the same trajectories.
@@ 6 Experiments
We conduct a comprehensive empirical evaluation of the Blind-IGT framework and the NLS algorithm. Our experiments aim to: (1) validate the theoretical convergence rates; (2) demonstrate the necessity of Blind-IGT when the temperature is unknown; and (3) assess robustness to model misspecification and unknown dynamics. (See Appendix F for detailed setup).
@@ 6.1 Experimental Setup
Markov Games. We simulate a zero-sum Markov Game with $|\mathcal{S}|=8$ states, $m=n=5$ actions, and discount factor $\gamma=0.9$. We adopt the Linear Q-function parameterization (Assumption 5.1) with $d=6$. The true temperature is set to $\tau^{*}=1.5$.
@@ 6.2 Validation of Convergence Rates
We validate the finite-sample guarantees established in Theorem 4.6 and Theorem 5.4.
@@
Markov Games Results. Figure 3 presents the results for the dynamic setting (assuming known P). The empirical slopes for the Q-parameters ($\hat{\theta}$) and the recovered reward function ($\hat{r}$) are -0.52, and -0.54 for temperature ($\hat{\tau}$). This strongly validates Theorem 5.4.
Figure 3: Convergence analysis in Markov Games (Theorem 5.4). The empirical slopes for Q-parameters ($\theta$), temperature ($\tau$), and the recovered reward function ($r$) closely match the theoretical $\mathcal{O}(K^{-1/2})$ rate.
Table 1: Comparison of Blind-IGT (NLS) vs. Standard IGT Liao et al. [2025] with known/misspecified temperature ($\tau^{*}=2.0$). Blind-IGT achieves near-oracle performance without prior knowledge of $\tau^{*}$.
@@ 6.3 The Necessity of Blind-IGT
We investigate the performance when the transition dynamics $P$ are unknown and must be estimated from data, relaxing the assumption made for Theorem 5.4.
We use the Markov Game setup ($|\mathcal{S}|=8,m=n=5,d=6$). We apply the NLS approach, estimating the transition dynamics $\hat{P}$ via Maximum Likelihood Estimation (with Laplace smoothing) from the observed trajectories. We compare this approach (Estimated-P) against an Oracle that uses the true dynamics (Known-P) during the reward recovery step.
@@ 7 Conclusion
This work establishes a new statistical foundation for Inverse Game Theory and multi-agent Inverse Reinforcement Learning by removing the restrictive assumption of known agent rationality. By rigorously addressing the bilinear coupling between rewards ($\theta$) and temperature ($\tau$), we resolved the fundamental scale ambiguity that previously precluded identification in entropy-regularized competitive games. Our analysis provides a complete characterization of identifiability through the statistically optimal NLS estimator, which achieves $\mathcal{O}(N^{-1/2})$ convergence rates.
Appendix A Algorithms
We present the full pseudocode for the algorithms proposed in this work.
Input: Data $\{(a^{k},b^{k})\}_{k=1}^{N}$, features $\phi$, Norm $C$
[…]
/* 1. Estimate QRE from observed frequencies */
[…]
/* 2. Construct regression matrices */
[…]
/* 3. Solve least squares problem */
$\hat{\theta}_{LS}\leftarrow(\hat{X}^{\top}\hat{X})^{-1}\hat{X}^{\top}\hat{y}$
/* 4. Recover temperature and reward parameters */
$\hat{\tau}\leftarrow C/\|\hat{\theta}_{LS}\|_{2}$; $\hat{\theta}\leftarrow\hat{\tau}\cdot\hat{\theta}_{LS}$
[…]
Algorithm 1 Normalized Least Squares (NLS) for Matrix Games
Input: Trajectories, features $\phi$, Norm $R$, Dynamics $P$ (or estimate)
[…]
/* 1. Estimate policies from trajectories */
$(\hat{\mu}(s),\hat{\nu}(s))\leftarrow$ frequency estimators for all $s\in\mathcal{S}$
/* 2. Estimate dynamics if unknown */
if $P$ unknown then
 $\hat{P}\leftarrow$ MLE from observed transitions
end if
/* 3. Construct aggregated system */
[…]
/* 4. Apply NLS procedure */
$\hat{\theta}_{LS}\leftarrow(\hat{X}^{\top}\hat{X})^{-1}\hat{X}^{\top}\hat{y}$; $\hat{\tau}\leftarrow R/\|\hat{\theta}_{LS}\|_{2}$; $\hat{\theta}\leftarrow\hat{\tau}\cdot\hat{\theta}_{LS}$
/* 5. Recover rewards via Bellman equation */
$\hat{Q}(s,a,b)\leftarrow\langle\phi(s,a,b),\hat{\theta}\rangle$ for all $(s,a,b)$
Compute $\hat{V}(s)$ from $\hat{Q},\hat{\mu},\hat{\nu},\hat{\tau}$
$\hat{r}(s,a,b)\leftarrow\hat{Q}(s,a,b)-\gamma\mathbb{E}_{s^{\prime}}[\hat{V}(s^{\prime})]$
[…]
Algorithm 2 Normalized Least Squares for Markov Games
Appendix B Notations
@@
$r,\,\tau,\,\gamma,\,P$ Reward function, temperature, discount factor, transition dynamics.
@@
$d,\,\phi$ Feature dimension and map. $L=\sup\|\phi\|_{2}$.
$\theta^{*},\,C,R$ True parameters; normalization constants for $\|\theta^{*}\|_{2}$.
@@
$\xi,\,C_{\min}$ Soft-Min Gap; coverage lower bound (Markov Games).
System Matrices
@@
$X(\mu,\nu),\,y(\mu,\nu)$ Aggregated feature matrix and log-ratio vector.
@@
$\hat{\theta}_{LS},\,\theta^{\prime*}$ LS solution (estimate of $\theta^{*}/\tau^{*}$); true unnormalized parameters.
Appendix C Background and Derivations